# Aeos: The Impact of Pop III Initial Mass Function and Star-by-Star Models in Galaxy Simulations

### v2

@@ 3.2 Chemical Enrichment
Nucleosynthetic yields from Pop III core-collapse SNe depend on their masses (Heger & Woosley, 2010). These yields differ not only in total metal production but also in the relative abundances of individual metals (Rossi et al., 2024; Brauer et al., 2025).
@@
Pop II stars in Aeos20 exhibit higher [X/Fe] values for light elements such as nitrogen (N) and oxygen (O), reflecting the increased nucleosynthetic yields from more massive Pop III stars due to stronger $\alpha$-capture reactions and rotational mixing (Heger & Woosley, 2010). Carbon (C) is also enhanced in Aeos20, primarily due to the significant production of carbon during helium burning in more massive Pop III stars. In these simulations we do not include pair-instability SNe (Rakavy & Shaviv, 1967; Fraley, 1968; Woosley et al., 2002), which may occur for high-mass ($\gtrsim 140$ M) Pop III stars (e.g. Schneider et al., 2004; Takahashi et al., 2016; de Bennassuti et al., 2017), but this would further enhance the amount of carbon due to their efficient ejection of carbon-rich material. For $\alpha$-elements such as magnesium (Mg) and calcium (Ca), Aeos20 also tends to show higher [X/Fe] ratios, consistent with the Heger & Woosley (2010) yields and the expectation that more massive Pop III stars are more efficient producers of $\alpha$-elements during their evolution and subsequent SN explosions.
@@
The enhanced carbon yield of more massive Pop III stars results in different amounts of carbon-enhanced metal-poor (CEMP) stars. CEMP stars, identified by their high carbon-to-iron ratios (taken either at [C/Fe] > 0.7 or [C/Fe] > 1.0), provide key insights into the formation of the first stars and galaxies (Beers & Christlieb, 2005; Jeon et al., 2017). Figure 6 shows the cumulative fractions of our simulated CEMP stars as a function of metallicity. Because more massive Pop III stars produce a higher ratio of carbon to iron, Aeos20 has a higher fraction of CEMP stars than Aeos10. While they reproduce the qualitative trends, compared with the observed fractions of CEMP stars in the Milky Way (Placco et al., 2014; Yoon et al., 2018), neither Aeos10 nor Aeos20 consistently reproduce observations CEMP stars at a given [Fe/H]. Generally, at the lowest metallicities that are most sensitive to Pop III enrichment, Aeos10 stars have lower fractions than observations and Aeos20 stars have higher fractions. This shows that the differences in Pop III initial mass functions between Aeos10 and Aeos20 significantly affect CEMP fractions and therefore points toward an improved IMF that can better reproduce Milky Way observations.
However, at the same time, there are inconsistencies in the fractions of carbon-enhanced metal-poor stars between different Milky Way stellar surveys (Arentsen et al., 2022; Ardern-Arentsen et al., 2025). Observations of dwarf galaxies also tend to find lower fractions of CEMP stars as compared to the Milky Way (e.g., Chiti et al. 2018; Sestito et al. 2024; Ou et al. 2025, Yelland et al. in prep). Furthermore, Pop III SNe likely exhibited a range of explosion energies (Kobayashi et al., 2006; Heger et al., 2003) that affected the metal retention of early galaxies (Cooke & Madau, 2014; Rossi et al., 2024). Our current simulations do not capture the effects of different explosion energies. Additional low-metallicity observations and simulations of larger boxes with a broader range of IMF variations and Pop III SNe explosion energies to further explore these discrepancies.
Figure 6: Cumulative fractions of carbon-enhanced metal-poor (CEMP) stars as a function of metallicity. Top panel shows CEMP stars with [C/Fe] > 0.7 and bottom panel shows CEMP stars with [C/Fe] > 1.0. The low-mass Aeos10 stars are shown as a teal line, the low-mass Aeos20 stars are shown as a maroon line, observations from Placco et al. (2014) are shown as diamonds, and observations from Yoon et al. (2018) are shown as circles. The Pop III IMF significantly affects the fraction of CEMP stars, and neither of the Aeos10 nor Aeos20 IMFs reproduce Milky Way observations.
@@
While there are differences in the enrichment of individual metals, the differences in Pop III IMF do not produce a noticeable difference in the stellar mass-metallicity relationships of the Aeos10 and Aeos20 simulations, as seen in Figure 7. This may be because the increased metal yields from more massive Pop III stars in Aeos20 are counterbalanced by their lower metal ejection efficiency. As modeled by Heger & Woosley (2010), massive Pop III stars often undergo fallback SNe, where a portion of the metals produced in the core region falls back onto the nascent black hole. This resulting metal retention depends on the explosion energy and stellar mass, and reduces the metals ejected into the interstellar medium and potentially limits the overall enrichment effect of these stars. Since we use the yields of Heger & Woosley (2010), our results inherently reflect this mechanism. Additionally, many of the more massive halos in our models have begun forming Pop II stars, which quickly dominate their hosts’ enrichment, obscuring differences from Pop III enrichment. To study the impact of different Pop III IMFs on chemical enrichment, it is therefore essential to focus exclusively on the oldest Pop II stars.
@@ 4 Comparison to Simulations Without Individual Stellar Feedback
Pop III stars are represented as individual particles in all simulations. However, the comparison simulations resolve Pop II stars as cluster particles of $\sim 1000$ M, as in Skinner & Wise (2020), while Aeos10 and Aeos20 resolve every Pop II star with $M>2$ M as an individual particle.
@@
The stellar masses, however, differ. The comparison simulations have a tendency to form galaxies of at least a few thousand solar masses because the Pop II star particles form in masses of $\sim$1000 M. The number of Pop II halos in each comparison simulation differs from just a few to a few dozen, but in aggregate, there is a trend towards more galaxies around 1000 M than in the Aeos simulations. The Aeos galaxies, on the other hand, can include Pop II galaxies with stellar masses as low as a few tens of solar masses (see Brauer et al., 2025). Additionally, both the Aeos simulations and comparison simulations can have stripped stars, resulting in the peak at a few M.
@@
The metallicities of the galaxies in the comparison simulations differ more significantly from those of the Aeos galaxies. Figure 7 shows the median mass-metallicity relation for each simulation with 16% to 84% scatter. The comparison simulations have a much shallower mass-metallicity relation than the Aeos simulations. The Aeos simulations have a log–log slope of $\sim 1$, which is steeper than the typical $\sim 0.3$ slope of other simulated mass-metallicity relations (Collins & Read, 2022), and much steeper than the practically flat relation of all the comparison simulations. The steeper Aeos slope in contrast to the slope of the comparison simulations may be due in part to the star-by-star modeling. Both our previous work and other star-by-star simulations (Brauer et al., 2025; Jeon & Ko, 2024; Andersson et al., 2025) have found that the relatively weaker feedback of the star-by-star method (compared to burstier, stronger feedback from traditional single stellar populations) leads to the possibility of higher stellar masses at a given halo mass – we start to see this in the stellar masses of our most massive galaxies, and likely do not see it at lower stellar masses in part due to the stellar mass resolution of the comparison simulations. This relatively weaker feedback could also allow greater retention of metals, and thus a steeper mass-metallicity relation as seen in Figure 7. In the comparison simulations, the aggregated bursty feedback from the more massive cluster particles (which quickly dominate galaxies with $M_{*}>10^{3}$ M) appear to eject metals more efficiently, reducing the expected increase in metallicity in halos between 103 M and 106 M in stellar mass. These halos are generally still low in total mass and have high loss fractions, making them highly sensitive to increases in feedback (see also Mead et al., 2025). We note, though, that the slope at higher halo mass is based on very few galaxies and the slope at lower halo mass is dominated by scatter. In particular, the difference in the largest galaxy and the metallicity of its surrounding galaxies between Aeos10 and the comparison galaxies is mostly due to a single burst of Pop II formation that occurs at the end of the Aeos10 run, so we refrain from drawing conclusions about $M_{*}>10^{5}$ M galaxies until longer-running simulations have been completed.
@@ 5 Conclusions
Our analysis of the Aeos simulations (Brauer et al., 2025) shows how varying the Pop III IMF and implementing a star-by-star feedback model influence early galaxy formation and evolution. We run two simulations with different Pop III IMFs: Aeos10, with a characteristic Pop III mass of 10 M and an upper cutoff of 100 M, and Aeos20, with a characteristic Pop III mass of 20 M and an upper cutoff of 300 M. By comparing the Aeos10 and Aeos20 simulations, we demonstrate that variations in the Pop III IMF significantly affect the ionization history, the number of early galaxies, and the chemical abundance patterns in subsequent Pop II stars. Aeos20, with its more massive Pop III stars, produces more ionizing photons, leading to substantially earlier ionization and consequently a suppression of small galaxy formation compared to what is occurring in Aeos10.
@@
We also run a suite of comparison simulations that lack individual star modeling. Comparisons with these simulations reveal that the high stellar resolution and star-by-star feedback in Aeos create differences in the stellar mass distribution and the mass-metallicity relation. While Pop II star formation is still relatively new, the comparison simulations tend to form more galaxies approximately around the mass of their Pop II aggregate star particle resolution, 1000 M, while the Aeos simulations avoid this by allowing Pop II stars to form from less massive gas reservoirs. Additionally, the star-by-star approach tends to result in galaxies with comparatively more stellar mass, as explored by Jeon & Ko (2024); Andersson et al. (2025); Brauer et al. (2025), due at least in part to less bursty feedback. This enables a greater accumulation of metals after Pop II star formation begins and consequently steeper mass-metallicity relations compared to simulations that represent stellar populations as single particles.

### v3

@@ 3.2 Chemical Enrichment
Nucleosynthetic yields from Pop III core-collapse SNe depend on their masses (HegerWoosley2010). These yields differ not only in total metal production but also in the relative abundances of individual metals (Rossi2024; AeosMethods).
@@
Pop II stars in Aeos20 exhibit higher [X/Fe] values for light elements such as nitrogen (N) and oxygen (O), reflecting the increased nucleosynthetic yields from more massive Pop III stars due to stronger $\alpha$-capture reactions and rotational mixing (HegerWoosley2010). Carbon (C) is also enhanced in Aeos20, primarily due to the significant production of carbon during helium burning in more massive Pop III stars. In these simulations we do not include pair-instability SNe (Rakavy1967; Fraley1968; Woosley2002), which may occur for high-mass ($\gtrsim 140$ M) Pop III stars (e.g. Schneider2004; Takahashi2016; deBennassuti2017), but this would further enhance the amount of carbon due to their efficient ejection of carbon-rich material. For $\alpha$-elements such as magnesium (Mg) and calcium (Ca), Aeos20 also tends to show higher [X/Fe] ratios, consistent with the HegerWoosley2010 yields and the expectation that more massive Pop III stars are more efficient producers of $\alpha$-elements during their evolution and subsequent SN explosions.
@@
The enhanced carbon yield of more massive Pop III stars results in different amounts of carbon-enhanced metal-poor (CEMP) stars. CEMP stars, identified by their high carbon-to-iron ratios (taken either at [C/Fe] > 0.7 or [C/Fe] > 1.0), provide key insights into the formation of the first stars and galaxies (Beers2005; Jeon2017). Figure 6 shows the cumulative fractions of our simulated CEMP stars as a function of metallicity. Because more massive Pop III stars produce a higher ratio of carbon to iron, Aeos20 has a higher fraction of CEMP stars than Aeos10. While they reproduce the qualitative trends, compared with the observed fractions of CEMP stars in the Milky Way (Placco2014; Yoon2018), neither Aeos10 nor Aeos20 consistently reproduce observations CEMP stars at a given [Fe/H]. Generally, at the lowest metallicities that are most sensitive to Pop III enrichment, Aeos10 stars have lower fractions than observations and Aeos20 stars have higher fractions. This shows that the differences in Pop III initial mass functions between Aeos10 and Aeos20 significantly affect CEMP fractions and therefore points toward an improved IMF that can better reproduce Milky Way observations.
However, at the same time, there are inconsistencies in the fractions of carbon-enhanced metal-poor stars between different Milky Way stellar surveys (Arentsen2022; Arentsen2025). Observations of dwarf galaxies also tend to find lower fractions of CEMP stars as compared to the Milky Way (e.g., Chiti18; Sestito2024; Ou2025, Yelland et al. in prep). Furthermore, Pop III SNe likely exhibited a range of explosion energies (Kobayashi2006; Heger2003) that affected the metal retention of early galaxies (Cooke2014; Rossi2024). Our current simulations do not capture the effects of different explosion energies. Additional low-metallicity observations and simulations of larger boxes with a broader range of IMF variations and Pop III SNe explosion energies to further explore these discrepancies.
Figure 6: Cumulative fractions of carbon-enhanced metal-poor (CEMP) stars as a function of metallicity. Top panel shows CEMP stars with [C/Fe] > 0.7 and bottom panel shows CEMP stars with [C/Fe] > 1.0. The low-mass Aeos10 stars are shown as a teal line, the low-mass Aeos20 stars are shown as a maroon line, observations from Placco2014 are shown as diamonds, and observations from Yoon2018 are shown as circles. The Pop III IMF significantly affects the fraction of CEMP stars, and neither of the Aeos10 nor Aeos20 IMFs reproduce Milky Way observations.
@@
While there are differences in the enrichment of individual metals, the differences in Pop III IMF do not produce a noticeable difference in the stellar mass-metallicity relationships of the Aeos10 and Aeos20 simulations, as seen in Figure 7. This may be because the increased metal yields from more massive Pop III stars in Aeos20 are counterbalanced by their lower metal ejection efficiency. As modeled by HegerWoosley2010, massive Pop III stars often undergo fallback SNe, where a portion of the metals produced in the core region falls back onto the nascent black hole. This resulting metal retention depends on the explosion energy and stellar mass, and reduces the metals ejected into the interstellar medium and potentially limits the overall enrichment effect of these stars. Since we use the yields of HegerWoosley2010, our results inherently reflect this mechanism. Additionally, many of the more massive halos in our models have begun forming Pop II stars, which quickly dominate their hosts’ enrichment, obscuring differences from Pop III enrichment. To study the impact of different Pop III IMFs on chemical enrichment, it is therefore essential to focus exclusively on the oldest Pop II stars.
@@ 4 Comparison to Simulations Without Individual Stellar Feedback
Pop III stars are represented as individual particles in all simulations. However, the comparison simulations resolve Pop II stars as cluster particles of $\sim 1000$ M, as in Skinner20, while Aeos10 and Aeos20 resolve every Pop II star with $M>2$ M as an individual particle.
@@
The stellar masses, however, differ. The comparison simulations have a tendency to form galaxies of at least a few thousand solar masses because the Pop II star particles form in masses of $\sim$1000 M. The number of Pop II halos in each comparison simulation differs from just a few to a few dozen, but in aggregate, there is a trend towards more galaxies around 1000 M than in the Aeos simulations. The Aeos galaxies, on the other hand, can include Pop II galaxies with stellar masses as low as a few tens of solar masses (see AeosMethods). Additionally, both the Aeos simulations and comparison simulations can have stripped stars, resulting in the peak at a few M.
@@
The metallicities of the galaxies in the comparison simulations differ more significantly from those of the Aeos galaxies. Figure 7 shows the median mass-metallicity relation for each simulation with 16% to 84% scatter. The comparison simulations have a much shallower mass-metallicity relation than the Aeos simulations. The Aeos simulations have a log–log slope of $\sim 1$, which is steeper than the typical $\sim 0.3$ slope of other simulated mass-metallicity relations (Collins2022), and much steeper than the practically flat relation of all the comparison simulations. The steeper Aeos slope in contrast to the slope of the comparison simulations may be due in part to the star-by-star modeling. Both our previous work and other star-by-star simulations (AeosMethods; Jeon2024; Andersson24) have found that the relatively weaker feedback of the star-by-star method (compared to burstier, stronger feedback from traditional single stellar populations) leads to the possibility of higher stellar masses at a given halo mass – we start to see this in the stellar masses of our most massive galaxies, and likely do not see it at lower stellar masses in part due to the stellar mass resolution of the comparison simulations. This relatively weaker feedback could also allow greater retention of metals, and thus a steeper mass-metallicity relation as seen in Figure 7. In the comparison simulations, the aggregated bursty feedback from the more massive cluster particles (which quickly dominate galaxies with $M_{*}>10^{3}$ M) appear to eject metals more efficiently, reducing the expected increase in metallicity in halos between 103 M and 106 M in stellar mass. These halos are generally still low in total mass and have high loss fractions, making them highly sensitive to increases in feedback (see also Mead2024b). We note, though, that the slope at higher halo mass is based on very few galaxies and the slope at lower halo mass is dominated by scatter. In particular, the difference in the largest galaxy and the metallicity of its surrounding galaxies between Aeos10 and the comparison galaxies is mostly due to a single burst of Pop II formation that occurs at the end of the Aeos10 run, so we refrain from drawing conclusions about $M_{*}>10^{5}$ M galaxies until longer-running simulations have been completed.
@@ 5 Conclusions
Our analysis of the Aeos simulations (AeosMethods) shows how varying the Pop III IMF and implementing a star-by-star feedback model influence early galaxy formation and evolution. We run two simulations with different Pop III IMFs: Aeos10, with a characteristic Pop III mass of 10 M and an upper cutoff of 100 M, and Aeos20, with a characteristic Pop III mass of 20 M and an upper cutoff of 300 M. By comparing the Aeos10 and Aeos20 simulations, we demonstrate that variations in the Pop III IMF significantly affect the ionization history, the number of early galaxies, and the chemical abundance patterns in subsequent Pop II stars. Aeos20, with its more massive Pop III stars, produces more ionizing photons, leading to substantially earlier ionization and consequently a suppression of small galaxy formation compared to what is occurring in Aeos10.
@@
We also run a suite of comparison simulations that lack individual star modeling. Comparisons with these simulations reveal that the high stellar resolution and star-by-star feedback in Aeos create differences in the stellar mass distribution and the mass-metallicity relation. While Pop II star formation is still relatively new, the comparison simulations tend to form more galaxies approximately around the mass of their Pop II aggregate star particle resolution, 1000 M, while the Aeos simulations avoid this by allowing Pop II stars to form from less massive gas reservoirs. Additionally, the star-by-star approach tends to result in galaxies with comparatively more stellar mass, as explored by Jeon2024; Andersson24; AeosMethods, due at least in part to less bursty feedback. This enables a greater accumulation of metals after Pop II star formation begins and consequently steeper mass-metallicity relations compared to simulations that represent stellar populations as single particles.